# The controlled exciton transport of the Multi-chain system by cavity-dressed energy level crossings and anticrossings

### v2

@@ II.3 Transport dynamics
where $\gamma_{\rm R}$ is the exciton exchanging rate between the last site of each chain and the drain, $\bar{n}_{\rm R}$ is the average exciton number in the drain. The last site of each chain exists coupling with the drain, so $k\in\{1,2,...,L\}$. In Eq. (16), the two terms denote respectively the reflection of excitons back to the last site of the $k$th chain and the exciton that transfer from the last column TLSs of the multichains to the drain. In order to show the net transportation of exciton, we set as $\bar{n}_{\rm R}=0$ which is reasonable as the photosynthesis process in nature. Therefore, the exciton is unidirectionally transferred from the last column of the multichains to the drain without exciton back.
@@ II.4 Efficiency of exciton transport
Figure 2: (color online) The energy spectra of the exciton-cavity coupling system as a function of the inter-chain coupling strength $\zeta$ for different configurations TC ST (a), SSH ST (b), TC-SSH ST (c), with the colormap representing the Hopfield coefficients of the content of cavity photons by solving the eigenvalues and eigenstates of the Hamiltonian (1) in zero- and single-excitation subspace. (d-f) The corresponding Hopfield coefficients of the content of cavity photons represented by line map. (g-i) The corresponding von Neumann entropy $S$ of excitons by calculating Eq. (13). The cavity-dressed energy level crossings (anticrossings) producing obvious effect on exciton transport are marked by the black (violet) vertical lines, which correspond to the saltatorial intersections (maximum values or acclive intersections) of the von Neumann entropy. The corresponding parameters are $L=2$, $N=8$, $\Delta_{\rm a}=\Delta_{\rm c}=0$ eV, $h_{\rm TC}=1$ eV, $h_{\rm SSH}=1$ eV, $\delta=0.5$, and $\Omega_{R}=1$ eV.
Figure 3: (color online) Similar as Fig 2, but for different configurations TC TT (a, d, g), SSH TT (b, e, h), and TC-SSH TT (c, f, i).
Figure 4: (color online) The energy spectra of the exciton-cavity coupling system to even chain $N=8$ as a function of the dimerization parameter $\delta$ for different configurations SSH ST (a), SSH TT (b), TC-SSH ST (c), and TC-SSH TT (d). The corresponding Hopfield coefficients of photons (e-h) and von Neumann entropy of excitons (i-l). The others are similar as the Fig. 2. The parameter used is $\zeta=0.1$ eV, the others are the same as those in Fig. 2.
@@
To show the exciton transport dynamics of the system containing a 2D multichains and a cavity, the exciton transport efficiency is defined and derived in this subsection. Generally, the exciton transport efficiency is characterized by the ratio of the outgoing $(I^{\rm s}_{\rm o})$ and input $(I^{\rm s}_{\rm i})$ exciton current for the steady state liu2019vibration ; wei2019enhanced , that is
@@
where the trace is operated overall the whole Hilbert space of the total system. The lefthand side of Eq. (19) means the rate of the change of occupation number with time, thus the first term (the second term) of the righthand side represents the input (outgoing) exciton current at the $j$th site in the $k$th chain. According to the explanation, under the TLS and TLS+cavity drive the input exciton current of the multichains can be written respectively as
@@
which is related to the exciton injection at the first site in the $k$th chain or/and in cavity via the external coherent field per unit time. Substituting Eq. (10) for $H_{\rm dc}$ and $H_{\rm dTc}$ in Eq. (20), under the zero- and single- excitation subspace, the input exciton current is simplified to
@@
According to the Hermitian property of density matrix, under the TLS and TLS+cavity drive the input exciton current is finally written as
@@
where $p_{k,N}$ is the occupation number at the last site in the $k$th chain.
@@ III Spectrum of coupling system
There are a large number of studies on the spectra of the TC and SSH Hamiltonian for the open boundary conditions campos2021generalization ; sirker2014boundary ; nie2020bandgap ; wei2022cavity ; cai2019observation . The TC chain can be regarded as a special configuration of the SSH chain. For the SSH chain, the topologically trivial and nontrivial phases depend on the parity of the number of an 1D TLSs chain and the qubit-qubit coupling configuration. According to the Ref. cai2019observation , the dimerization parameter $\delta>0$ ($\delta<0$) corresponds to the topological winding number $\nu=1$ ($\nu=0$), i.e., the topologically nontrivial (trivial) regime. In Ref. sirker2014boundary , the results show that there is one zero-energy mode in whole range of $\delta$ and are two localized nearly-zero-energy modes in the topologically nontrivial regime $\delta>0$ for the spectra of odd and even chain respectively. The finite gap between the two edge modes leads to the formation of the hybrid edge modes that overlap with each other, generating the polaritonic states when the edge modes are dressed by cavity due to the coupling between the excitons and the photons. Comparing to the single chain system, the presence of the inter-chain coupling induces the extension of the spectra of the 2D multi-chain system. As we all known, the gap corresponds to the energy level anticrossings, at which the LZ transitions among the levels occur in multi-level LZ system, adding greatly to the complexity of the system dynamics hu2018effect . What’s more, the energy level crossings mean that the corresponding quantum states are degeneracy. In the exciton-cavity coupling system, at the crossings the excitations are coherently exchanged between the TLSs and the cavity. Therefore, the crossings and anticrossings of the edge modes dressed by cavity can lead to the abundant exchange of photon between the multichains and the cavity, increasing or decreasing polaritons. The forming polaritons act as intermediary in exciton transport, where the excitons are transported among site $(k,1)~{}\rightarrow$ cavity $\rightarrow$ site $(k,N)$ without experiencing the dissipation of the other TLSs on multichains wei2022cavity . Therefore, the generating of the polaritons leads to the exciton transport efficiency and currents increasing, which can be controlled by the crossings and anticrossings.
@@
Figure 6: (color online) The dependence of the steady-state occupation numbers (a1 and b1) $p_{11}$ on site 11, (a2 and b2) $p_{\rm bulk}$ on bulk sites, (a3 and b3) input exciton current $I_{\rm i}$, (a4 and b4) outgoing exciton current $I_{\rm o}$, and (a5 and b5) transport efficiency $\eta$ on the inter-chain coupling strength $\zeta$ (a1-5) and the dimerization parameter $\delta$ (b1-5) for different configurations without cavity. In (b1-5) the solid and dash line correspond respectively to $N=8$ and 9. The parameters used are $\zeta=0.1$ eV, $\xi_{{\rm T},1}=\xi_{\rm c}=0.1$ eV, $\gamma_{\rm d}=0.01$ eV, $\gamma_{\rm R}=0.1$ eV, and the others are the same as those in Fig. 2.
@@
Figure 7: (color online) The dependence of the steady-state dynamics of the hybrid system on the inter-chain coupling strength $\zeta$ in the presence of cavity under TLS drive (a1-6) and TLS+cavity drive (a1-6). (a5-b5) The occupation number of photon with $\zeta$, in which the inflection points are marked by the vertical fine lines, and the others are similar to Fig. 6. The parameter used is $\kappa=0.01$ eV, and the others are the same as those in Fig. 6.
@@
In order to analyse the spectra of the coupling system consisting of the multichains and the cavity, its eigenvalues can be obtained by numerically diagonalizing the Hamiltonian (1). On the other hand, the formation of polaritons can be investigated by calculating the Hopfield coefficients $|\alpha_{k,i;c}|^{2}$. When the $j$th normalized eigenvector is expanded as $|\psi_{j}\rangle=\Sigma_{k}^{L}\Sigma_{i}^{N}\alpha_{k,i}|k,i\rangle+\alpha_{c}|c\rangle$, $|\alpha_{k,i;c}|^{2}$ denotes Hopfield coefficients, which characterize the corresponding hybridizing weight of different excitonic and photonic contents. Hence, the color maps Fig. 2 (a-c), 3 (a-c) and Fig. 4 (a-d), 5 (a-d) show respectively the spectra of the exciton-cavity coupling system as functions of the inter-chain coupling strength $\zeta$ and the dimerization parameter $\delta$ for different coupling configurations, containing the photonic Hopfield coefficients $|\alpha_{c}|^{2}$. To show obviously the photonic Hopfield coefficients, we plot the line maps of $|\alpha_{c}|^{2}$ of each eigenstate as functions of $\zeta$ and $\delta$ in Fig. 2 (d-f), 3 (d-f) and Fig. 4 (e-i), 5 (e-i) respectively. To find out the crossings and anticrossings in spectra, the corresponding von Neumann entropy $S$ of excitons is calculated through Eq. (13), and is shown in Fig. 2 (g-i), 3 (g-i), 4 (i-l), and 5 (i-l). One can see that the zero-energy modes exist for each situation. It is obvious that only the nearly-zero-energy modes are distinctly dressed by the cavity mode, generating corresponding polaritonic branches, because the high excited states are hardly dressed for the given coupling strength between excitons and photons. At the crossings (anticrossings) one can see that the photonic contents exchange between the upper and lower polaritonic branches with $\zeta$ and $\delta$, but the distributed ratio between the upper and lower polaritonic branches depends on the locations of the crossings (anticrossings). It means that the exchange of excitation or the LZ transition of photon occurs, inducing the change of the exciton transport efficiency and currents, which will be illustrated in the next section by the steady-state dynamics. Further, it can be discovered that the effect of the inter-chain coupling strength on the crossings (anticrossings) is more significant than that of the dimerization parameter. The number of crossings (anticrossings) affecting transport is relate to the coupling configuration in the range of $\zeta$. Comparing to the changes of $|\alpha_{c}|^{2}$ and of $S$ with $\zeta$ and $\delta$, a common phenomenon can be seen that the locations of the intersections of the photonic Hopfield coefficients and of the von Neumann entropies corresponding to different eigenstates are consistent.
@@
In the next section, the steady-state dynamics with $\zeta$ and $\delta$ will be shown, in which the relation between the dressed crossings (anticrossings) and the occupation number of photons can be analysed. The changes of exciton transport currents and efficiency can be explained. We aim at revealing the mechanism of exciton transport by finding the optimal range of parameters on exciton transport and comparing the spectra and steady-state dynamics for different configurations.
@@
Figure 9: (color online) The dependence of the steady-state dynamics on the inter-chain coupling strength $\zeta$ and the detuning $\Delta$ of the driving frequency to the TLS without cavity under TLS drive (a-f, ab-fb), and with cavity under TLS drive (g-l, gc-lc) and TLS+cavity (the TLS and cavity drive exist simultaneously) drive (m-r) for different configurations. The corresponding physical quantities of each row are shown at right. The zones that cavity produces significant effect on exciton transport efficiency are marked by ellipses in (g-l), the corresponding zone are marked in (ab-fb) and (gc-lc). The parameters used are the same as those in Fig. 7.
@@
Figure 11: (color online) The steady-state dynamics as a function of the number $L$ of the chains without (a1111-g1111) and with (a2222-h2222) cavity for different configurations, in which $p_{1,i}$, $p_{L,i}$, and $p_{bulk-in}$ are respectively represent the occupation number of the first row, the last row, and the bulk state denoted as ”bulk-in” containing all TLSs inside of multi-chain system. The parameter used is $N=10$, and the others are the same as those in Fig. 7.
Figure 12: (color online) The steady-state dynamics as a function of the length $N$ of chain for different configurations and number of the chains, (a-d) without cavity under TLS drive, (e-h) with cavity under TLS drive. The 1,2,3,4th column correspond to $L=2,3,4,5$ respectively. The parameters used are the same as those in Fig. 7.

### v3

@@ II.3 Transport dynamics
where $\gamma_{\rm R}$ is the exciton exchanging rate between the last site of each chain and the drain, $\bar{n}_{\rm R}$ is the average exciton number in the drain. The last sites of each chain exist coupling with the drain, so $k\in\{1,2,...,L\}$. In Eq. (15), the two terms denote respectively the inflow and outflow of excitons from the last column TLSs of the multichains to the drain. In order to show the net transportation of exciton, we set as $\bar{n}_{\rm R}=0$ which is reasonable as the photosynthesis process in nature. Therefore, the exciton is unidirectionally transferred from the last column of the multichains to the drain without exciton reflection.
@@ II.4 Efficiency of exciton transport
Figure 2: (color online) The energy spectra of the exciton-cavity coupling system as a function of the inter-chain coupling strength $\zeta$ for different configurations HO ST (a), HE ST (b), HO-HE ST (c), with the colormap representing the Hopfield coefficients of the content of cavity photons by solving the eigenvalues and eigenstates of the Hamiltonian (1) in zero- and single-excitation subspace. (d-f) The second order partial derivative of the Hopfield coefficients of the content of cavity photons with respect to $\zeta$, which contain the discontinuity and the extremum corresponding to the marked locations existing evident changes in steady-state dynamics. (g-i) The corresponding von Neumann entropy $S$ of excitons by calculating Eq. (12). The color of the lines has not means just to distinguish each other in (d-i). The cavity-dressed energy level crossings (anticrossings) producing obvious effect on exciton transport are marked by the black (violet) vertical lines and the data texts of their locations, which correspond to the saltatorial intersections (maximum values or acclive intersections) of the von Neumann entropy. The corresponding parameters are $L=2$, $N=8$, $\Delta_{\rm a}=\Delta_{\rm c}=0$ eV, $h=1$ eV, $\delta=0.5$, and $\Omega_{R}=1$ eV.
Figure 3: (color online) Similar as Fig 2, but for different configurations HO TT (a, d, and g), HE TT (b, e, and h), and HO-HE TT (c, f, and i).
Figure 4: (color online) Similar as Fig 2, but for even chain $N=8$ and for different configurations HE ST (a, e, and i), HE TT (b, f, and j), HO-HE ST (c, g, and k), and HO-HE TT (d, h, and l) with dimerization parameter $\delta$. The parameter used is $\zeta=0.1$ eV, the others are the same as those in Fig. 2.
@@
To show the exciton transport dynamics of the system containing a 2D multichains and a cavity, the exciton transport efficiency is defined and derived in this subsection. Generally, the exciton transport efficiency is characterized by the ratio of the outgoing $(I^{\rm s}_{\rm o})$ and input $(I^{\rm s}_{\rm i})$ exciton currents for steady state liu2019vibration ; wei2019enhanced , that is
@@
where the trace is operated throughout the Hilbert space of the total system. The lefthand side of Eq. (18) means the rate of the change of $p_{k,j}(t)$ with time, thus the first term (the second term) of the righthand side represents the input (outgoing) exciton current at the $j$th site in the $k$th chain. According to the explanation, under the TLS and TLS+cavity drive the input exciton current of the multichains can be written respectively as
@@
which is related to the exciton injection at the first site in the $k$th chain or/and in cavity via the external coherent field per unit time. Substituting Eq. (9) for $H_{\rm dc}$ and $H_{\rm dTc}$ in Eq. (19), under the zero- and single- excitation subspace, the input exciton currents are simplified to
@@
According to the Hermitian property of density matrix, under the TLS and TLS+cavity drive the input exciton currents are finally written as
@@
where $p_{k,N}$ is the occupation number on the last site of the $k$th chain.
@@ III Spectrum of coupling system
There are a large number of studies on the spectra of the SSH Hamiltonian for the open boundary conditions sirker2014boundary ; nie2020bandgap ; wei2022cavity ; cai2019observation . The topological phases of the SSH chain depend on the parity of the number of an 1D TLSs chain and the qubit-qubit coupling configuration. According to the Ref. cai2019observation , the dimerization parameter $\delta>0$ ($\delta<0$) corresponds to the topological winding number $\nu=1$ ($\nu=0$), i.e., the topologically nontrivial (trivial) regime. In Ref. sirker2014boundary , the spectra show that there is one zero-energy mode in whole range of $\delta$ for odd and are two localized nearly-zero-energy modes in the topologically nontrivial regime $\delta>0$ for even chain. The finite gap between the two edge modes leads to the formation of the hybrid edge modes that overlap with each other, generating the polaritonic states when the edge modes are dressed by cavity due to the coupling between the excitons and the photons. Comparing to the single chain system, the presence of the inter-chain coupling induces the extension of the spectra of the 2D multi-chain system. As we all known, the gap corresponds to the energy level anticrossings, at which the LZ transitions among the levels occur in multi-level LZ system, adding greatly to the complexity of the system dynamics hu2018effect . What’s more, the energy level crossings mean that the corresponding quantum states are degeneracy. In the exciton-cavity coupling system, at the crossings the excitations are coherently exchanged between the TLSs and the cavity. Therefore, the crossings and anticrossings of the edge modes dressed by the cavity can lead to the abundant exchange of excitation between the multichains and the cavity, increasing or decreasing polaritons. The forming polaritons act as intermediary in exciton transport, where the excitons are transported among site $(k,1)~{}\rightarrow$ cavity $\rightarrow$ site $(k,N)$ without experiencing the dissipation of the other TLSs on multichains wei2022cavity . Therefore, the polaritons lead to the exciton transport efficiency and currents increasing, which can be controlled by the crossings and anticrossings.
@@
Figure 6: (color online) The dependence of the steady-state occupation numbers (a1 and b1) $p_{11}$ on site 11, (a2 and b2) $p_{\rm bulk}$ on bulk sites, (a3 and b3) input exciton current $I_{\rm i}$, (a4 and b4) outgoing exciton current $I_{\rm o}$, and (a5 and b5) transport efficiency $\eta$ on the inter-chain coupling strength $\zeta$ (a1-5) and the dimerization parameter $\delta$ (b1-5) for different configurations without cavity. In (b1-5) the solid and dash lines correspond respectively to $N=8$ and 9. The parameters used are $\zeta=0.1$ eV, $\xi_{{\rm T},1}=\xi_{\rm c}=0.1$ eV, $\gamma_{\rm d}=0.01$ eV, $\gamma_{\rm R}=0.1$ eV, and the others are the same as those in Fig. 2.
@@
Figure 7: (color online) The dependence of the steady-state dynamics of the hybrid system on the inter-chain coupling strength $\zeta$ in the presence of cavity under TLS drive (a1-6) and TLS+cavity drive (b1-6). (a5 and b5) The occupation number of photon with $\zeta$, in which the extreme points are marked by the vertical fine lines, and the others are similar to Fig. 6. The parameter used is $\kappa=0.01$ eV, and the others are the same as those in Fig. 6.
@@
In order to analyse the spectra of the coupling system consisting of the multichains and the cavity, its eigenvalues can be obtained by numerically diagonalizing the Hamiltonian (1). Further, the formation of polaritons can be investigated by calculating the Hopfield coefficients $|\alpha_{c}|^{2}$. When the $j$th normalized eigenvector is expanded as $|\psi_{j}\rangle=\Sigma_{k}^{L}\Sigma_{i}^{N}\alpha_{k,i}|k,i\rangle+\alpha_{c}|c\rangle$, $|\alpha_{k,i;c}|^{2}$ denotes Hopfield coefficient, which characterizes the corresponding hybridizing weight of different excitonic and photonic contents. The color maps Fig. 2 (a-c), 3 (a-c) and Fig. 4 (a-d), 5 (a-d) show respectively the spectra of the exciton-cavity coupling system as functions of the inter-chain coupling strength $\zeta$ and the dimerization parameter $\delta$ for different coupling configurations, and contain the photonic Hopfield coefficient $|\alpha_{c}|^{2}$. One can see that the zero-energy modes exist in the spectra for each situation. $|\alpha_{c}|^{2}\neq 0$ means that the energy level is dressed by the cavity mode. It is obvious that only the nearly-zero-energy modes are distinctly dressed by the cavity mode, generating corresponding polaritonic branches, because the high excited states are hardly dressed for the given coupling strength between excitons and photons. In addition, there are the crossings (anticrossings) in the spectra, in which the crossings (anticrossings) between the nearly-zero-energy modes dressed by the cavity mode are our focus. It can be seen that the ratio of the distributed photonic contents on the upper and lower polaritonic branches vary with $\zeta$ and $\delta$. It means that the coherent exchange of excitation or the LZ transition of photons occurs, inducing the change of the exciton transport efficiency and currents, which will be analysed in the next section by the steady-state dynamics. Further, it can be discovered that the number and locations of crossings (anticrossings) affecting transport depend on the coupling configurations of TLSs in the range of $\zeta$ and $\delta$. In comparison, the effect of $\zeta$ on the crossings (anticrossings) is more significant than that of $\delta$.
In order to find the changing regularity of the photonic Hopfield coefficient $|\alpha_{c}|^{2}$, we plot the scatter maps of the second order partial derivatives $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ of $|\alpha_{c}|^{2}$ with respect to $\zeta$ in Fig. 2 (d-f), 3 (d-f) and $\partial^{2}|\alpha_{c}|^{2}/\partial\delta^{2}$ with respect to $\delta$ in Fig. 4 (e-i), 5 (e-i) that are related to the obvious changes of the steady-state dynamics of exciton transport. To determine and distinguish accurately the locations of the crossings and anticrossings in spectra, the corresponding von Neumann entropy $S$ of excitons is calculated through Eq. (12), and is shown in Fig. 2 (g-i), 3 (g-i), 4 (i-l), and 5 (i-l). The crossings (anticrossings) determined by $S$ is signed in the figure of the spectra, which correspond to the saltatorial intersections (maximum values or acclive intersections) of $S$. Significantly, we discovery that the incontinuity locations of $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ and $\partial^{2}|\alpha_{c}|^{2}/\partial\delta^{2}$ with $\zeta$ and $\delta$ correspond to the crossings, and the extreme points of that correspond to the anticrossings. As shown in Fig. 2(d), the green and red lines of $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ are discontinuity at $\zeta=-0.3$ where the dressed energy levels exist the crossing [Fig. 2(a) and 2(g)], and the green and cyan lines of that appear the maximal values as well as the blue and green lines of that appear the minimal values at $\zeta=-0.21$, -0.02 respectively where the dressed energy levels exist the anticrossings [Fig. 2(a) and 2(g)]. We also can see that the green, blue, and cyan lines are discontinuity at $\zeta=0.07$, 0.09, 0.22 respectively [Fig. 2(e)] corresponding to the crossings of the dressed energy levels [Fig. 2(b) and 2(h)]. The green and magenta lines behave as maximal values, and the magenta and cyan lines behave as minimal values at $\zeta=-0.03$, 0.48 respectively [Fig. 2(e)] corresponding to the anticrossings of the dressed energy levels [Fig. 2(b) and 2(h)]. Similarly, the maximal values occur at $\zeta=-0.1$, -0.05, and 0.03 in the black, green, and blue lines respectively, and the minimal values occur at $\zeta=-0.05$, 0.03 in the red and cyan lines respectively [Fig. 2(f)], which correspond to the anticrossings of the dressed energy levels [Fig. 2(c) and 2(i)]. Therefore, the second order partial derivative of the photonic Hopfield coefficients is a good physical quantity to determine the crossings and anticrossings of spectra. Comparing the spectra under different configurations, one can discover that the gap in the range $\zeta\in[-0.21,-0.02]$ for the configuration HO ST and its range [Fig. 2(a)] is respectively widened and extended by the heterogeneous coupling in multichain as shown in Fig. 2(b). For the mixed configuration HO-HE ST new gaps appear near $\zeta=0$ as shown in Fig. 2(c). About $\partial^{2}|\alpha_{c}|^{2}/\partial\zeta^{2}$ and $\partial^{2}|\alpha_{c}|^{2}/\partial\delta^{2}$ in Fig. 3 (d-f), 4 (e-i), and 5 (e-i), the conclusion is similar as that of Fig. 2 (d-f), thus it is no longer stated hereinbelow. When the inter-chain coupling is TT, the spectra is shown in Fig. 3(a-c). It can be seen that the gaps near $\zeta=0$ narrow, and its ranges in $\zeta$ reduce compared with Fig. 2(a-c). In the space of $\delta$, the spectra are presented in Fig. 4 (a-d) for even chain and Fig. 5 (a-d) for odd chain. In Fig. 4 (a-d) we can see that the configuration HE ST has more crossings of and energy levels of the dressed nearly-zero-energy modes than HE TT. In mixed multichains the presence of the HO chain leads also to the decrease of the crossings and energy levels, comparing Fig. 4 (a) and 4 (c). However, the crossings at $\delta=0.3$ and 0.36 [Fig. 4 (c)] covert to anticrossing at $\delta=0.19$ [Fig. 4 (d)] because of TT coupling. The energy levels of the dressed nearly-zero-energy modes cross for even chain and avoid crossing for odd chain. Therefore, there is less crossings affecting exciton transport for odd chain [Fig. 5 (a-d)].
@@
In the next section, the steady-state dynamics with $\zeta$ and $\delta$ will be shown, in which the relation between the dressed crossings (anticrossings) and the occupation number of photons can be discussed. The changes of exciton transport currents and efficiency can be explained. We aim at revealing the mechanism of exciton transport by finding the optimal range of parameters on exciton transport and comparing the spectra and steady-state dynamics for different configurations.
@@
Figure 9: (color online) The dependence of the steady-state dynamics on the inter-chain coupling strength $\zeta$ and the detuning $\Delta$ of the driving frequency relative to the TLS without cavity under TLS drive (a-f, ab-fb), and with cavity under TLS drive (g-l, gc-lc) and under TLS+cavity (the TLS and cavity drive exist simultaneously) drive (m-r) for different configurations. The corresponding physical quantities of each row are shown at right. The zones that cavity produces significant effect on exciton transport efficiency are marked by ellipses in (g-l), the corresponding zones are marked in (ab-fb) and (gc-lc). The parameters used are the same as those in Fig. 7.
@@
Figure 11: (color online) The steady-state dynamics as a function of the number $L$ of the chains without a1-g1) and with (a2-h2) cavity for different configurations, in which $p_{1,i}$, $p_{L,i}$, and $p_{\rm bulk-in}$ represent respectively the occupation number of the first row, the last row, and the bulk state denoted as ”bulk-in” containing all TLSs inside of the multichains. The parameter used is $N=10$, and the others are the same as those in Fig. 7.
Figure 12: (color online) The steady-state dynamics as a function of the length $N$ of chains for different configurations and number of the chains, (a-d) without cavity under TLS drive, (e-h) with cavity under TLS drive. The parameters used are the same as those in Fig. 7.